OpenStudy (nir1232):

For the reaction CH4 + 2O2 → CO2 + 2H2O, how many moles of carbon dioxide are produced from the combustion of 221.6 g of methane? Select one: a. 13.81 b. 27.61 c. 3557 d. 4.612

OpenStudy (idahochik):

Ok so do you know how to do stoichiometry? Because thats how you solve this problem

OpenStudy (idahochik):

you start by setting it up like this \[given *\frac{ information about what is wanted }{ information about given }\]

OpenStudy (idahochik):

so for your equation it would look like this\[221.6 CH _{4}\times \frac{ 12.011gCO _{2} }{ 16.04gCH _{4} }\]

OpenStudy (idahochik):

the CH4s cross out and when you do all the arithmetic you come out with 165.9gCO2

OpenStudy (idahochik):

To convert that to moles you set it up like this \[165.9gCO _{2}\times \frac{ 1 mole }{ 12.011gCO _{2} }\]

OpenStudy (idahochik):

when you divide all of that out you come out with 13.81 moles of CO2 which would make your answer a. hope this helped!

OpenStudy (sweetburger):

What @idahochik is perfectly fine. However, I would have done it like this \[221.6g(CH_4)\times \frac{ 1 mole(CH_4) }{ 16.04g(CH_4) }\] This would give you the moles of CH4. Looking at the molar coefficients of the given balanced reaction we know that the ratio of CH4 to CO2 is 1:1. Therefore the number of moles of CH4 that are used is equivalent to the number of moles of CO2 produce.